Solve the systems by graphing
3x-2y=-24
-3x-y=6
I put them in point slope for
I think:
y=3/2x + 12
y=-3x-6
Then created a T-chart
x y
0 12
2 15
x y
0 -6
1 -9
Can you please help me? I do not understand why I am stuck....

Answer by ewatrrr(24785)


Hi
3x-2y=-24 OR y= 3/2x + 12 (Green)
x=0, then y= 12
y=0 then x =-8
(0,12)and(-8,0)
-3x-y=6 OR y=-3x-6 (Blue)
x=0, then y=-6
y=0 then x= -2
(0,-6) and (-2,0)
Plotting points and connecting with Line
Lines Intersect at ordered pair Pt(-4,6)
This is the solution for this system